OpenStudy (3psilon):

How to simplify ((1/x)-(1/3)) / (x-3) ?

OpenStudy (3psilon):

\[\frac{ \frac{ 1 }{ x } - \frac{ 1 }{ 3 } }{ x-3 }\]

OpenStudy (anonymous):

-1/3x

OpenStudy (3psilon):

But can I see the steps @jk_16

OpenStudy (anonymous):

here is the idea: (a/b)-(c/d) = (-b c+a d)/(b d)

OpenStudy (anonymous):

do that for the top part..then do it again in the step when you divide by x-3

OpenStudy (3psilon):

Why divide that by (bd)?

OpenStudy (3psilon):

@jk_16

OpenStudy (anonymous):

you multiply the denominator

OpenStudy (3psilon):

the denominator is (x-3) though why (bd) ?

OpenStudy (anonymous):

you have a choice you can multiply top and bottom by \(3x\) to clear the fractions, or you can actually do the addition in the numerator and then simplify the compound fraction

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ x } - \frac{ 1 }{ 3 } }{ x-3 }=\frac{\frac{3-x}{3x}}{x-3}=\frac{3-x}{3x}\times \frac{1}{x-3}\] \[=\frac{1}{3x}\times \frac{3-x}{x-3}\] and since \(\frac{3-x}{x-3}=-1\) you are left with \(-\frac{1}{3x}\)

OpenStudy (3psilon):

How is it equal to -1? @satellite73